A stock has a beta of 1.19 and an expected return of 10.43 percent. If the stock's reward-to-risk ratio is 6.68 percent, what is the risk-free rate?
Answer #1

SOLUTION:

Reward-to-risk ratio = (Expected return - Risk-free rate ) / Beta

0.0668 = (0.1043- Risk-free rate) / 1.19

(0.1043 - Risk-free rate) = 0.0668 × 1.19

(0.1043 - Risk-free rate) = 0.079492

Risk-free rate = 0.1043 - 0.079492 = 2.48%

Risk-free rate = 0.0265 or 2.65%